# Block Vecchia Approximation for Scalable and Efficient Gaussian Process Computations

### v1

@@ 3 Numerical Studies
In this section, we evaluate the accuracy and computational performance of our proposed block Vecchia algorithm compared to the classical Vecchia algorithm. Our block Vecchia code was compiled using gcc version 10.2.0 (12.2.0) and CUDA version 11.4 (11.8). It was linked with Intel MKL 2022.2.1, MAGMA 2.6.0 (2.7.2), GSL 2.6, and NLopt v2.7.1 optimization libraries. All computations were performed in double-precision arithmetic, and we conducted each experimental run five times to verify repeatability (time to solution). We employed the deterministic metric (KL divergence) for accuracy assessment and qualitative analysis and conducted simulation studies under various parameter configurations. We also compare the block Vecchia method with exact GPs in terms of parameter estimation and predictive performance in two real-world applications.
@@ 3.1 KL Divergence
In this subsection, we calculate the deterministic metric, KL divergence, to assess the accuracy of the block Vecchia algorithm with the help of ExaGeoStat (Abdulah et al.,, 2018a), on Gaussian random fields with $n=20,000$ spatial locations within $[0,1]\times[0,1]$. The KL divergence is defined as:
@@
We investigate the impact of permutation and block count of the block Vecchia algorithm. As illustrated in Figure 5a, the permutation plays an essential role in enhancing the accuracy of the block Vecchia algorithm, where the maxmin permutation (mmd in figures) achieves the highest accuracy, random permutation yields near-optimal accuracy, while other permutations fail to produce promising results. In Figure 5b, we set random reordering as the default, recognizing that it achieves near-optimal accuracy while demanding fewer computational resources compared to maxmin reordering (Guinness,, 2018). The KL divergence is plotted against different block counts, consistently confirming that a larger block count leads to more accurate results. This enhancement can be attributed to two factors: firstly, a smaller block size (or a larger block count) ensures that the neighbors are representative of the points within the block; secondly, random ordering potentially yields better neighbor candidates for blocks. Figure 5 only presents the case of $\beta=0.052537$ and $\nu=1.5$. Refer to Figure S1 for more parameters settings.
@@
In our analysis of two approximation methodologies, the classic Vecchia exhibits a pronounced decrease in accuracy as smoothness increases, while the block Vecchia demonstrates considerable robustness, maintaining accuracy across a range of smoothness parameters. For the evaluation of the block Vecchia approximation capabilities concerning smoothness parameters, we focus on the case where $bc=1500$, identified as both relatively efficient and accurate. Figure 6 provides quantitative evidence of this phenomenon: for the CV-20K-150-random configuration, KL divergence escalates from $10^{-4}$ to $10^{-1}$, indicating a substantial loss in approximation precision. Conversely, the BV-1500-450-random configuration shows only a slight decrease in KL divergence, from $10^{-4.6}$ to $10^{-4}$, highlighting the effectiveness of the BV method to manage the complexities associated with high smoothness levels effectively. Furthermore, as $\nu$ increases, the gap between the block Vecchia and classic Vecchia methods diminishes. Refer to Figure S3 for more parameters settings.
@@
The block Vecchia algorithm outperforms the classic Vecchia algorithm in both computational efficiency and accuracy. To evaluate the efficiency and accuracy of the block Vecchia, we focus on the case at $bc=1500$, which is the same as above. In Figure 7, the time-to-solution is defined as the cumulative duration required to compute the log-likelihood for a single iteration in MLE. This includes the time for matrix generation and batched BLAS operations while explicitly excluding the time spent on nearest neighbor searching and K-means clustering. These latter operations are omitted from consideration due to their one-time execution at the outset of MLE and their minimal impact on the overall computation time across numerous iterations. The BV-1500-:-random achieves equivalent levels of accuracy approximately 5X faster than its classic counterpart, CV-20K-:-random. Additionally, it offers a more precise KL approximation. Notably, the block Vecchia provides a more accurate approximation within the same time frame compared to the classic Vecchia, CV-20K-:-random. Refer to Figure S4 for more parameters settings.
@@ 3.2 Simulations for Parameter Estimation and Prediction
In this section, we assess the statistical parameter estimation accuracy and the prediction uncertainty of the block Vecchia. Using the ExaGeoStat framework (Abdulah et al.,, 2018a), a high-performance unified framework for computational geostatistics on many-core systems, we generate 50 datasets on irregular 2D spatial locations within the unit square $[0,1]\times[0,1]$ to investigate the parameter uncertainty numerically, which has been studied for exact GP asymptotically (Wang et al.,, 2023). These datasets are based on Gaussian random fields with a problem size of $n=20,000$ observations and the Matérn covariance function, which adheres to the parameter configuration outlined in Table 1. We estimate the parameters set $(\sigma^{2},\beta,\nu)$ using the BOBYQA optimization algorithm (Powell et al.,, 2009), which is a gradient-free method and has great power to search the global optimal value. For our approximation methodologies, we select the classic Vecchia approach with a conditioning size of 30 as our baseline, following the recommendation made by (Guinness,, 2018; Pan et al.,, 2024). This conditioning size is also chosen based on its demonstrated capability to provide comparable approximation accuracy in terms of KL divergence, as evidenced in Figure 6 for the classic Vecchia algorithm. In contrast, for the block Vecchia approximation, we set the block count, $bc=1500$, identified as offering a balance between accuracy and computational efficiency relative to the classic approach, with varying conditioning sizes of $(10,30,60,120,180)$ explored.
@@
3.3 BV performance on large-scale real datasets.
@@ 4 Application to 3D Wind Speed Profiles
Traditional GP models are computationally infeasible for such large datasets, particularly when handling the vertical dimension in 3D profiles. However, the block Vecchia approximation, by breaking down the computations into smaller blocks, enables the processing of large-scale 3D wind profiles, capturing fine vertical and horizontal wind patterns with much greater efficiency and scalability. This scalability is crucial for analyzing vast datasets across multiple altitudes and spatial regions, which were previously inaccessible due to computational constraints. In this section, we focus on modeling high-resolution 3D wind speed profiles. The scalability of the block Vecchia method is evaluated on a single GPU (NVIDIA V100 with 32 GB memory) with problem sizes at the million level. The 17 atmospheric layers of the 3D profile are included as an additional input dimension in the 3D profile wind speed data, e.g., Figure S10 illustrates the residuals. The wind speed measurements across the 17 layers span 37 million irregular locations, posing a significant computational challenge for modern computers. We employ a subsampling technique to address this, randomly select 1 million locations and scale the coordinates to $(x,y,z)\in[0,1]^{3}$. The rest of the experimental configuration remains consistent with the previous section, and we then apply the block Vecchia approximation to estimate the parameters.
@@
Figure 13 presents the results of parameter estimation for modeling the residuals of 3D wind speed in a million-level context and the MSPE. The number of blocks increases (e.g., 10K, 50K, and 100K), reflecting our interest in improving accuracy in the approximation of the block Vecchia method. The findings indicate that: 1) parameter estimation gradually converges to a specific value as the conditioning size increases; 2) the convergence rate improves with a larger block count. The experiment encourages extending the Vecchia approximation to larger-scale problems. These results demonstrate that larger conditioning sizes and block counts enhance the accuracy of parameter estimation. More importantly, the block Vecchia method can handle much larger problem sizes than the classic Vecchia, facilitating large-scale modeling. We also assessed the predictive ability of the block Vecchia method. We plug in the converged estimated parameters $\hat{\theta}$ into the block Vecchia approximation, then conduct 1000 rounds of the conditional simulations, and finally report the MSPE and the standard deviation. Specifically, Figure 13 (d) illustrates the MSPE of the block Vecchia approximation for the residuals of 3D wind speed. We have used different block counts (BV-5000, BV-10000, BV-20000, BV-30000) and varied the conditioning size for prediction in the approximated GP framework, where the block Vecchia method is applied for scalable approximations. It is observed that 1) increasing the conditioning size improves the prediction accuracy, and using a larger block count ($10{,}000$-$30{,}000$) yields more stable results even at smaller conditioning sizes; 2) the gains in prediction accuracy increases at a certain level of conditioning size for all types of block counts.
@@ 5 Conclusion
This study introduces the block Vecchia algorithm and its GPU framework based on batched operations provided by the MAGMA linear algebra library. The algorithm evaluates in batch, simultaneously processing multiple operations or data points, and simultaneously computes the multivariate conditional likelihood of all location blocks to improve efficiency, reduce storage requirements, and enhance scalability in large-scale scenarios. Our numerical study and real datasets analysis provide a deep insight into the block Vecchia algorithm: 1) The analysis reveals that a larger conditioning size and block count can improve modeling accuracy compared to the classic Vecchia algorithm, ensuring efficiency and precision; 2) The sequence of blocks plays a critical role in the accuracy of the approximation, i.e., the random ordering markedly enhances the approximation as the number of blocks increases; 3) The block Vecchia method demonstrates an approximately $80$X speedup compared to the classic Vecchia algorithm without compromising the accuracy of the approximations. This significant enhancement in computational speed, coupled with improved accuracy, represents a substantial advancement in applying Gaussian process models; and 4) scalability: The block Vecchia method allows handling problem sizes $40$X larger than those accommodated by the classic Vecchia algorithm. This scalability enables the algorithm to leverage existing GPUs effectively, making it a powerful tool for addressing large-scale statistical modeling challenges. According to the asymptotic properties of Vecchia approximation (Kang et al.,, 2024; Zhang et al.,, 2021), the asymptotic properties of the MLEs from the block Vecchia method will be investigated in our future research. Besides, considering the promising results of the classic Vecchia approximation on high-dimension problems Katzfuss et al., (2022); Jimenez and Katzfuss, (2023), it is worth studying the computer experiment with the block Vecchia.
@@
S3 Smoothness enhancement
We investigate the numerical accuracy of block Vecchia algorithm with regards to parameters in the Matérn kernel. Figure S3 indicates that the block Vecchia method improves KL divergence as smoothness increases, narrowing the gap between the block Vecchia and classic Vecchia and achieving higher overall accuracy with larger conditioning set.
@@
S4 Faster and more accurate numerical results
We investigate the time and accuracy of the block Vecchia algorithm with regards to parameters in the Matérn kernel. Figure S4 indicates that the block Vecchia method achieve more accurate KL divergence within a shorter time.
@@
S5 Numerical results of increasing number of locations

### v2

@@ 3 Numerical Studies
This section assesses the accuracy and computational performance of the proposed block Vecchia algorithm compared to the classical Vecchia algorithm. The block Vecchia code was compiled using GCC version 10.2.0 (or 12.2.0) and CUDA v11.4 (or v11.8), and linked with Intel MKL v2022.2.1, MAGMA v2.6.0 (or v2.7.2), GSL v2.6, and NLopt v2.7.1 optimization libraries. Our machine has 40 CPU cores (Intel(R) Xeon(R) Gold 6230R CPU @ 2.10GHz) and NVIDIA V100 GPU 32GB. Each experiment was repeated five times to ensure repeatability and consistency in the time-to-solution metric. Accuracy was evaluated using the deterministic KL divergence metric alongside qualitative analysis, with simulation studies conducted under various parameter settings. Furthermore, the block Vecchia algorithm was compared with exact GPs in terms of parameter estimation and predictive performance across two real-world applications.
@@ 3.1 KL Divergence
In this subsection, we calculate the deterministic metric, KL divergence, to assess the accuracy of the block Vecchia algorithm with the help of ExaGeoStat (Abdulah et al.,, 2018a), on Gaussian random fields with $n=20000$ spatial locations within $[0,1]\times[0,1]$. We also include additional small-scale examples in S8 and S9 for consumer-grade GPUs, using a dataset size of $n=8000$.
The KL divergence is defined as:
@@
We investigate the impact of permutation and block count of the block Vecchia algorithm. As illustrated in Figure 5a, the permutation plays an essential role in enhancing the accuracy of the block Vecchia algorithm, where the maxmin permutation (mmd in figures) achieves the highest accuracy, random permutation yields near-optimal accuracy, while other permutations fail to produce promising results. In Figure 5b, we set random reordering as the default, recognizing that it achieves near-optimal accuracy while demanding fewer computational resources than the maxmin reordering algorithm (Guinness,, 2018). The KL divergence is plotted against different block counts, consistently confirming that a larger block count leads to more accurate results. This enhancement can be attributed to two factors: firstly, a smaller block size (or a larger block count) ensures that the neighbors are representative of the points within the block; secondly, random ordering potentially yields better neighbor candidates for blocks. Figure 5 only presents the case of $\beta=0.052537$ and $\nu=1.5$. Refer to Figure S1 for more parameter settings.
@@
In our analysis of two approximation methodologies, the classic Vecchia exhibits a pronounced decrease in accuracy as smoothness increases. At the same time, the block Vecchia demonstrates considerable robustness, maintaining accuracy across a range of smoothness parameters. For the evaluation of the block Vecchia approximation capabilities concerning smoothness parameters, we focus on the case where $bc=1500$, identified as both relatively efficient and accurate. Figure 6 provides quantitative evidence of this phenomenon: for the CV-20K-150-random configuration, KL divergence escalates from $10^{-4}$ to $10^{-1}$, indicating a substantial loss in approximation precision. Conversely, the BV-1500-450-random configuration shows only a slight decrease in KL divergence, from $10^{-4.6}$ to $10^{-4}$, highlighting the effectiveness of the BV method to manage the complexities associated with high smoothness levels effectively. Furthermore, as $\nu$ increases, the gap between the block Vecchia and classic Vecchia methods diminishes. Refer to Figure S3 for more parameter settings.
@@
The block Vecchia algorithm outperforms the classic Vecchia algorithm in both computational efficiency and accuracy. To evaluate the efficiency and accuracy of the block Vecchia, we focus on the case at $bc=1500$, which is the same as above. In Figure 7, the time-to-solution is defined as the cumulative duration required to compute the log-likelihood for a single iteration in MLE. This includes the time for matrix generation and batched BLAS operations while explicitly excluding the time spent on nearest neighbor searching and K-means clustering. These latter operations are omitted from consideration due to their one-time execution at the outset of MLE and their minimal impact on the overall computation time across numerous iterations. The BV-1500-:-random achieves equivalent levels of accuracy approximately 5X faster than its classic counterpart, CV-20K-:-random. Additionally, it offers a more precise KL approximation. Notably, the block Vecchia provides a more accurate approximation within the same time frame compared to the classic Vecchia, CV-20K-:-random. Refer to Figure S4 for more parameter settings. In addition, we provide the accuracy of block Vecchia approximation along with the increasing number of locations in S5.
@@ 3.2 Simulations for Parameter Estimation and Prediction
In this section, we assess the accuracy of statistical parameter estimation and the prediction uncertainty of the block Vecchia. Using the ExaGeoStat framework (Abdulah et al.,, 2018a), a high-performance unified framework for computational geostatistics on many-core systems, we generate 50 datasets on irregular 2D spatial locations within the unit square $[0,1]\times[0,1]$ to investigate the parameter uncertainty numerically, which has been studied for exact GP asymptotically (Wang et al.,, 2023). These datasets are based on Gaussian random fields with a problem size of $n=20,000$ observations and the Matérn covariance function, which adheres to the parameter configuration outlined in Table 1. We estimate the parameters set $(\sigma^{2},\beta,\nu)$ using the BOBYQA optimization algorithm (Powell et al.,, 2009), which is a gradient-free method and has great power to search the global optimal value. For our approximation methodologies, we select the classic Vecchia approach with a conditioning size of 30 as our baseline, following the recommendation made by (Guinness,, 2018; Pan et al.,, 2024). This conditioning size is also chosen based on its demonstrated capability to provide comparable approximation accuracy in terms of KL divergence, as evidenced in Figure 6 for the classic Vecchia algorithm. In contrast, for the block Vecchia approximation, we set the block count, $bc=1500$, identified as offering a balance between accuracy and computational efficiency relative to the classic approach, with varying conditioning sizes of $(10,30,60,120,180)$ explored.
@@
3.3 Block Vecchia Performance on Large-Scale Real Datasets
@@ 4 Application to 3D Wind Speed Profiles
Traditional GP models are computationally infeasible for such large datasets, particularly when handling the vertical dimension in 3D profiles. However, by breaking down the computations into smaller blocks, the block Vecchia approximation enables the processing of large-scale 3D wind profiles, capturing fine vertical and horizontal wind patterns with much greater efficiency and scalability. This scalability is crucial for analyzing vast datasets across multiple altitudes and spatial regions, which were previously inaccessible due to computational constraints. In this section, we focus on modeling high-resolution 3D wind speed profiles. The scalability of the block Vecchia method is evaluated on a single GPU (NVIDIA V100 with 32 GB memory) with problem sizes at the million level. The 17 atmospheric layers of the 3D profile are included as an additional input dimension in the 3D profile wind speed data, e.g., Figure S10 illustrates the residuals. The wind speed measurements across the 17 layers span 37 million irregular locations, posing a significant computational challenge for modern computers. We employ a subsampling technique to address this, randomly select 1 million locations and scale the coordinates to $(x,y,z)\in[0,1]^{3}$. The rest of the experimental configuration remains consistent with the previous section, and we then apply the block Vecchia approximation to estimate the parameters.
@@
Figure 13 presents the results of parameter estimation for modeling the residuals of 3D wind speed in a million-level context and the MSPE. The number of blocks increases (e.g., 10K, 50K, and 100K), reflecting our interest in improving accuracy in the approximation of the block Vecchia method. The findings indicate that: 1) parameter estimation gradually converges to a specific value as the conditioning size increases; 2) the convergence rate improves with a larger block count. The experiment encourages extending the Vecchia approximation to larger-scale problems. These results demonstrate that larger conditioning sizes and block counts enhance the accuracy of parameter estimation. More importantly, the block Vecchia method can handle much larger problem sizes than the classic Vecchia, facilitating large-scale modeling. We also assessed the predictive ability of the block Vecchia method. We plug in the converged estimated parameters $\hat{\theta}$ into the block Vecchia approximation, then conduct 1000 rounds of the conditional simulations, and finally report the MSPE and the standard deviation. Specifically, Figure 13 (d) illustrates the MSPE of the block Vecchia approximation for the residuals of 3D wind speed. We have used different block counts (BV-5000, BV-10000, BV-20000, BV-30000) and varied the conditioning size for prediction in the approximated GP framework, where the block Vecchia method is applied for scalable approximations. It is observed that 1) increasing the conditioning size improves the prediction accuracy, and using a larger block count ($10{,}000$-$30{,}000$) yields more stable results even at smaller conditioning sizes; 2) the gains in prediction accuracy increases at a certain level of conditioning size for all types of block counts. In addition, we provide a full dataset application on soil moisture with 2M points in S7.
@@ 5 Conclusion
This study introduces the block Vecchia algorithm and its GPU framework based on batched operations provided by the MAGMA linear algebra library. The algorithm evaluates in batch, simultaneously processing multiple operations or data points, and simultaneously computes the multivariate conditional likelihood of all location blocks to improve efficiency, reduce storage requirements, and enhance scalability in large-scale scenarios. Our numerical study and real datasets analysis provide a deep insight into the block Vecchia algorithm: 1) The analysis reveals that a larger conditioning size and block count can improve modeling accuracy compared to the classic Vecchia algorithm, ensuring efficiency and precision; 2) The sequence of blocks plays a critical role in the accuracy of the approximation, i.e., the random ordering markedly enhances the approximation as the number of blocks increases; 3) The block Vecchia method demonstrates an approximately $80$X speedup compared to the classic Vecchia algorithm without compromising the accuracy of the approximations. This significant enhancement in computational speed, coupled with improved accuracy, represents a substantial advancement in applying Gaussian process models; and 4) scalability: The block Vecchia method allows handling problem sizes $40$X larger than those accommodated by the classic Vecchia algorithm. This scalability enables the algorithm to leverage existing GPUs effectively, making it a powerful tool for addressing large-scale statistical modeling challenges. In a recent study (Hazra et al.,, 2024), the Vecchia approximation (represented by the GpGp package (Guinness et al.,, 2021)) was shown to outperform many popular methods for fitting Gaussian Processes across five large spatial datasets under various parameter settings. This confirms the superiority of the classic Vecchia approach. Accordingly, we applied our proposed block Vecchia algorithm to all datasets from this study, evaluating its performance against the results reported in the study and the local approximation Gaussian Process (laGP, (Gramacy,, 2016)). Our results, presented in Table S1 in the supplementary material S10, demonstrate that the block Vecchia algorithm is as accurate as GpGp/GpGp0, while our GPU implementation achieves an average of $1.5$X speedup in modeling the five given datasets. According to the asymptotic properties of Vecchia approximation (Kang et al.,, 2024; Zhang et al.,, 2021), the asymptotic properties of the MLEs from the block Vecchia method will be investigated in our future research. Besides, considering the promising results of the classic Vecchia approximation on high-dimension problems (Katzfuss et al.,, 2022; Jimenez and Katzfuss,, 2023), it is worth studying computer experiments with block Vecchia.
@@
6 Acknowledgment
This research was supported by King Abdullah University of Science and Technology (KAUST). We gratefully acknowledge the funding and resources provided by KAUST, which made this work possible and has facilitated the research and analysis presented in this study.
@@ 6 Acknowledgment
S3 Accuracy of Block Vecchia at Varying Smoothness Levels
We investigate the numerical accuracy of the block Vecchia algorithm using Matérn kernel. Figure S3 shows that the block Vecchia method improves KL divergence as smoothness increases, narrowing the gap between the block Vecchia and classic Vecchia and achieving higher overall accuracy with larger conditioning set.
@@
S4 Additional Results on Block Vecchia Accuracy and Time
We analyze the time efficiency and accuracy of the block Vecchia algorithm in relation to the parameters of the Matérn kernel. Figure S4 demonstrates that the block Vecchia method achieves a more accurate KL divergence in a shorter time.
@@
S5 Impact of Increasing the Number of Locations
@@
S8 Numerical Study with Increased Block Count: A Small-Scale Study
We investigate the numerical accuracy of the block Vecchia algorithm, focusing on the impact of block count and different reorderings in a smaller problem, $n=8000$. Figure S14 illustrates the KL divergence as the block count increases across three different configurations, with fixed parameters $\beta=0.052537$ and $\nu=1.5$. Our findings indicate that the maxmin reordering (mmd in figures) achieves the highest accuracy, random reordering yields near-optimal accuracy, while other reorderings fail to produce promising results. Additionally, a higher block count improves the accuracy of the block Vecchia method for both random and maxmin reorderings. In Figure S15, we use random reordering as the default, recognizing that it achieves near-optimal accuracy while demanding fewer computational resources than the maxmin reordering. Across various parameter settings, the KL divergence is plotted against different block counts, consistently confirming that a larger block count leads to more accurate results.
(a) $\beta=0.052537,bc=200$
(b) $\beta=0.052537,bc=1500$
(c) $\beta=0.052537,bc=2500$
(a) $\beta=0.026270,\nu=0.5$
(b) $\beta=0.078809,\nu=0.5$
(c) $\beta=0.210158,\nu=0.5$
(d) $\beta=0.017512,\nu=1.5$
(e) $\beta=0.052537,\nu=1.5$
(f) $\beta=0.140098,\nu=1.5$
(g) $\beta=0.014290,\nu=2.5$
(h) $\beta=0.042869,\nu=2.5$
(i) $\beta=0.114318,\nu=2.5$
S9 Accuracy of Block Vecchia at Varying Smoothness Levels: A Small-Scale Study
We investigate the numerical accuracy of the block Vecchia algorithm in terms of parameters in the Matérn kernel at $n=8000$. Figure S16 indicates that the block Vecchia improves KL divergence as smoothness increases, narrowing the gap between the block Vecchia and classic Vecchia and achieving higher overall accuracy with larger conditioning set.
(a) $\beta=0.026270,\nu=0.5$
(b) $\beta=0.078809,\nu=0.5$
(c) $\beta=0.210158,\nu=0.5$
(d) $\beta=0.017512,\nu=1.5$
(e) $\beta=0.052537,\nu=1.5$
(f) $\beta=0.140098,\nu=1.5$
(g) $\beta=0.014290,\nu=2.5$
(h) $\beta=0.042869,\nu=2.5$
(i) $\beta=0.114318,\nu=2.5$
Figure S16: KL divergence and conditioning size under 8K locations with log10 scale under $bc=650$.
S10 Experiment on Competition Spatial Dataset
Hazra et al., (2024) explored various statistical and deep learning methods to address the challenges of analyzing large spatial datasets. They compared the following popular methods: Vecchia approximation, Block Composite Likelihood (BCL), fixed rank kriging (FRK), Gapfill, local approximate Gaussian Process (laGP), Stochastic Partial Differential Equation (SPDE), Nearest Neighbor Gaussian Process (NNGP), and deep learning approaches, e.g., DeepGP and DeepKriging. Their results demonstrated that statistical methods outperformed other approaches across various datasets, particularly BCL, the Vecchia approximation (GpGp), and its zero-mean variant (GpGp0). These methods achieved lower prediction errors and narrower uncertainty intervals. We added the proposed BV approximation algorithm to the comparison using only one dataset from this study (dataset 3). We have compared the BV and laGP to the existing methods in Hazra et al., (2024). Specifically, we set the block count as $3000$, the conditioning size as $270$ for the block Vecchia, and the total size of the local designs in the laGP as $60$, the same as the conditioning size in GpGP0. The results are summarized in the last two columns of Table S1, with the results of the first nine models copied from Hazra et al., (2024). The block Vecchia approximation is observed to be faster and achieve accuracy comparable to GpGp0, GpGp, and BCL, with these four methods yielding the best accuracy.
Table S1: Model comparison (Hazra et al.,, 2024) for the five datasets of Sub-competition 2a (Hong et al.,, 2023) on different metrics.
Models BCL FRK Gapfill GpGp GpGp0 SPDE NNGP DeepKriging DeepGP laGP BV
Dataset 1
MSPE 0.274 0.274 0.282 0.274 0.274 0.274 0.278 0.318 0.317 0.277 0.274
MAPE 0.417 0.417 0.424 0.417 0.417 0.417 0.420 0.449 0.448 0.420 0.417
PICP 0.951 0.951 0.964 0.953 0.953 0.951 0.950 0.939 0.961 0.942 0.951
MPIW 2.049 2.050 2.258 2.075 2.074 2.050 2.058 2.144 2.310 2.023 2.056
Time (s) 372 209 8 547 595 10 618 392 2197 158 462
Dataset 2
MSPE 0.274 0.276 0.347 0.275 0.275 0.276 0.279 0.313 0.297 0.278 0.274
MAPE 0.417 0.419 0.470 0.418 0.418 0.419 0.421 0.441 0.421 0.420 0.417
PICP 0.951 0.950 0.983 0.953 0.952 0.951 0.952 0.942 0.961 0.944 0.951
MPIW 2.053 2.058 3.246 2.076 2.076 2.058 2.076 2.159 2.330 2.040 2.059
Time (s) 403 156 8 510 478 8 568 393 1951 159 377
Dataset 3
MSPE 0.297 0.489 0.983 0.297 0.297 0.419 0.298 0.380 0.829 0.337 0.297
MAPE 0.434 0.557 0.788 0.434 0.434 0.515 0.436 0.490 0.721 0.462 0.434
PICP 0.951 0.951 0.974 0.952 0.952 0.949 0.954 0.951 0.963 0.946 0.951
MPIW 2.137 2.744 4.391 2.147 2.147 2.543 2.177 2.418 3.830 2.265 2.143
Time (s) 555 129 8 514 521 9 626 381 2233 163 380
Dataset 4
MSPE 0.275 0.277 0.346 0.275 0.275 0.277 0.279 0.319 0.296 0.278 0.275
MAPE 0.418 0.420 0.468 0.418 0.418 0.419 0.421 0.447 0.433 0.420 0.418
PICP 0.951 0.951 0.967 0.953 0.953 0.951 0.951 0.951 0.955 0.944 0.951
MPIW 2.054 2.062 2.556 2.077 2.077 2.060 2.071 2.141 2.200 2.040 2.062
Time (s) 475 145 8 528 494 8 630 379 1759 173 273
Dataset 5
MSPE 0.302 0.540 1.004 0.302 0.302 0.455 0.304 0.409 0.956 0.358 0.302
MAPE 0.438 0.586 0.797 0.438 0.438 0.537 0.440 0.509 0.779 0.477 0.438
PICP 0.951 0.950 0.976 0.951 0.952 0.951 0.955 0.962 0.964 0.943 0.951
MPIW 2.154 2.885 4.576 2.162 2.162 2.648 2.206 2.493 4.088 2.316 2.160
Time (s) 340 116 8 565 585 7 648 380 1630 169 323
Code Archive
The Block Vecchia algorithm provides a Github repository, https://github.com/paper-code1/BV-Gaussian. You can follow the README.md to install the package and reproduce the results in the paper.